OpenStudy (anonymous):

Needs to chose two animals in a race. Zebra runs 40mph and elephant runs 25mph. Converted it to yards per second. 19.56 yards/second and 12.22 yards/second respectively. Second part is stumping me. Decide how much of a head start the faster animal offers the slower animal (in yards). Write equations for each animal relating distance from the starting line to time. I then need to graph them on the same coordinate plane and determine how long (in seconds and in yards) it will take the faster to overtake the slower.

OpenStudy (unklerhaukus):

\[s_z=19\frac59[\text{yd/s}]\qquad s_e=12\frac29[\text{yd/s}]\] \[s=\frac dt\] \[t=\frac ds\] for a good race the expected time for each contestant is the same \[t_z=t_e\] \[\frac{d_z}{s_z}=\frac {d_e}{s_e}\] \[\frac{s_e}{s_z}=\frac{d_e}{d_z}\]